# Counting rational maps on $\mathbb{P}^{1}$ with prescribed local conditions

### v2

@@ 1.1. Background and motivation
For an elliptic curve defined over a number field, one can define the local minimal discriminant at each non-archimedean place and define the (global) minimal discriminant as the product of the local minimal discriminants [Sil09, Chapters VII–VIII]. Similarly, for a rational map defined over a number field, one can define the local minimal resultant at each non-archimedean place and define the (global) minimal resultant as their product, see [Sil07, Chapter 4.11]. After homogenization, a rational map $\phi\in\mathbb{Q}(z)$ is given by a pair of homogeneous polynomials $F(X,Y),G(X,Y)\in\mathbb{Z}[X,Y]$ without common factors in $\mathbb{Z}[X,Y]$. We say that $\phi$ is globally minimal or has globally minimal resultant if the resultant of $F$ and $G$ is equal to the minimal resultant of $\phi$. Over a complete algebraically closed non-archimedean valued field, the notion of (local) minimal resultants give rise to the so called minimal resultant locus which has interesting analytic and moduli-theoretic characterizations [Rum15, Rum17]. Over $\mathbb{Q}$, globally minimal maps play an important role in a conjecture of Silverman on the uniformity of number of integral points in orbits. This is, in turn, analogous to a conjecture of Lang for integral points on elliptic curves after work of Dem’janenko, see [BM12] and the reference therein for more details.
In recent years, there have been various results concerning statistical aspects of the arithmetic dynamics on $\mathbb{P}^{1}$, for example [HH19, Hin21a, Hin21b, DHHJ24]. However, the nature of these results involves a stochastic process in which one successively composes a random rational map from a finite set endowed with a probability measure. On the other hand, our work continues the classical analogy program between arithmetic dynamics on $\mathbb{P}^{1}$ and elliptic curves. In arithmetic statistics of elliptic curves, considerable interest lies in questions regarding the distribution of sets $\mathcal{S}$ with certain properties. For instance, the rank distribution conjecture of Katz and Sarnak predicts that exactly $1/2$ (resp. $1/2$) of the elliptic curves over $\mathbb{Q}$ have rank 00 (resp. $1$). Cremona and Sadek [CS23] consider the density of global Weierstrass models of elliptic curves over $\mathbb{Q}$ with prescribed reduction types at the primes. We prove analogues of these results for models for dynamical systems, arising from rational maps of prescribed degree $d\geq 2$ in $\mathbb{P}^{1}$.
@@ 1.2. Main results
$P(\phi):=[A_{0}:\cdots:A_{d}:B_{0}:\cdots:B_{d}]\in\mathbb{P}^{2d+1}(\mathbb{Q}).$
The height of $\phi$, denoted $H(\phi)$, is defined to be the absolute multiplicative Weil height of the point $P(\phi)$ [BG06, Chapter 1]. For any $x>0$, let $\mathcal{R}^{(d)}(x)=\{\phi\in\mathcal{R}^{(d)}:\ H(\phi)\leq x\}$. It is seen that
#$\#\mathcal{R}^{(d)}(x)=\frac{(2x)^{2d+2}}{2\zeta(2d+2)}+O\left(x^{2d+1}\log x\right),$
@@
$\liminf_{x\rightarrow\infty}\left(\frac{\#\{\phi\in\mathcal{R}^{(d)}(x)\mid\phi\in\mathcal{S}\}}{\#\mathcal{R}^{(d)}(x)}\right)>c.$
Inspired by Poonen’s work [Poo03], there is a more general density to consider. First, it is possible to regard $\mathcal{R}^{(d)}\subset\mathbb{P}^{1}(\mathbb{Q})$ as subsets of $\mathbb{Z}^{2d+2}$, see Notation 2.2. Then one may consider the box density in [Poo03, p. 355]. There are weaker versions of this called the weak box densities, cf. Theorem 3.2 of loc. cit. More specifically, for $\mathcal{S}\subset\mathcal{R}^{(d)}$ and for each natural number $n\in[1,2d+2]$, we have the notion of the weak box density $\mu_{n}(\mathcal{S})$, see Definition 2.11. The following analogue of the result of Cremona-Sadek [CS23] arises: with respect to the box density, does a positive proportion of rational maps have globally minimal resultant? This question, along with related topics, will be explored in a future work.
@@ 1.3. Methodology
We briefly discuss the techniques used in the proof of Theorems A, B and C. A classical result of Schanuel [Sch79] gives the asymptotic count for the number of rational points in projective space with height $\leq x$. For two generic homogeneous polynomials $F(X,Y)$ and $G(X,Y)$ of degree $d$, their resultant is a homogeneous polynomial in the coefficients of $F$ and $G$. Hence $F$ and $G$ have a common factor if and only if the resultant polynomial vanishes. A result by Schwartz and Zippel (cf. Lemma 2.5) provides an upper bound on the number of zeros of this polynomial with height $\leq x$. Combining this result with Schanuel’s result implies that
#$\#\mathcal{R}^{(d)}(x)\sim\frac{(2x)^{2d+2}}{2\zeta(2d+2)},$
as detailed in Proposition 2.6. Following Poonen’s work [Poo03], we can also study the “box version” of the above results.
In Section 3, we study properties of subsets $\mathcal{S}\subseteq\mathcal{R}^{(d)}$ defined by the so called local conditions. We define local densities at each prime $p$. Proposition 3.4 employs a sieve-theoretic argument to show that, if a certain admissibility condition is met, the density of $\mathcal{S}$ is the product of local densities at all primes $p$. These results are applied in Section 4 to prove Theorems A and B. Although these local conditions are insufficient to describe the entire collection of $\phi$ with minimal resultant, they provide positive lower bounds on the density. This relies on explicit local density calculations, with a main innovation involving counting the number of $\mathbb{Z}/p^{2}\mathbb{Z}$-points on the integral scheme defined by the resultant polynomial and its partial derivatives. Some similar arguments have been employed in [Poo03]. In Section 5, an explicit primary decomposition for this integral scheme is used to obtain an explicit lower bound on the weak box densities when $d=2$. Our computations are aided by Macaulay2.
@@ 2. The height density, box density, and weak box densities
We define the height $H(\phi)$ of $\phi\in\mathcal{R}^{(d)}$ as the absolute multiplicative Weil height of the corresponding point in $\mathbb{P}^{2d+1}(\mathbb{Q})$. Then we define the height density of a subset $\mathcal{S}$ of $\mathcal{R}^{(d)}$. Using classical results of Schanuel [Sch79] for the distribution of rational points on projective spaces and the Schwart-Zipple lemma [Sch80, Zip79], we derive an asymptotic formula for
$\mathcal{R}^{(d)}(x):=\{\phi\in\mathcal{R}^{(d)}:\ H(\phi)\leq x\}.$
We also recall the notion of box density and weak box densities for subsets of $\mathbb{Z}^{2d+2}$ from Poonen’s work [Poo03]. We have the analogues of Schanuel’s result and an asymptotic formula for counting $\mathcal{R}^{(d)}$ in a box; however the statement of these analogues is delayed until the next section since the proof relies on certain sieving results. Finally we define the box density and weak box densities for a subset $\mathcal{S}$ of $\mathcal{R}^{(d)}$.
@@ 2.1. The height density
$P=P(\phi):=[A_{0}:\cdots:A_{d}:B_{0}:\cdots:B_{d}]\in\mathbb{P}^{2d+1}(\mathbb{Q}).$
@@
[…]
@@
Thus we count models for dynamical systems, and not isomorphism classes. Nonetheless, this seems to be natural way to order rational maps, the definition is analogous to counting Weierstrass equations for elliptic curves according to their height (cf. [CS23]). For a detailed treatment of heights, we refer to [BG06].
@@
We shall also denote the coordinates of $Q$ by $(A_{0},\dots,A_{d},B_{0},\dots,B_{d})$, i.e., $A_{i}:=C_{i+1}$ and $B_{j}:=C_{d+j+2}$. Note that if $Q$ satisfies (2), then exactly one of the points $Q$ or $-Q$ satisfies (1).
@@
$H(\phi)=\max\{|A_{i}|,|B_{j}|:0\leq i,j\leq d\},$
see [BG06, Chapter 1].
@@
$\mathcal{S}(x)=\{P\in\mathcal{S}:\ H(P)\leq x\}.$
@@
(2.2) #$\#\mathbb{P}^{2d+1}(\mathbb{Q},x)=\frac{1}{2\zeta(2d+2)}(2x)^{2d+2}+O\left(x^{2d+1}\log x\right).$
@@
This is from [Sch79, Theorem 1]. ∎
@@
$\{(x_{1},\ldots,x_{n})\in I_{1}\times\cdots\times I_{n}:\ f(x_{1},\ldots,x_{n})=0\}$
@@
This was proven independently by Schwartz [Sch80] and Zippel [Zip79]. ∎
@@
#$\#\mathcal{R}^{(d)}(x)=\frac{1}{2\zeta(2d+2)}(2x)^{2d+2}+O\left(x^{2d+1}\log x\right),$
@@
#$\#(\mathbb{P}^{2d+1}(\mathbb{Q},x)\setminus\mathcal{R}^{(d)}(x))=O(x^{2d+1})$
@@
$\mathfrak{d}(\mathcal{S}):=\lim_{x\rightarrow\infty}\frac{\#\mathcal{S}(x)}{\#\mathbb{P}^{2d+1}(\mathbb{Q},x)},$
@@
$\underline{\mathfrak{d}}(\mathcal{S}):=\liminf_{x\rightarrow\infty}\frac{\#\mathcal{S}(x)}{\#\mathbb{P}^{2d+1}(\mathbb{Q},x)}\text{ and }\overline{\mathfrak{d}}(\mathcal{S}):=\limsup_{x\rightarrow\infty}\frac{\#\mathcal{S}(x)}{\#\mathbb{P}^{2d+1}(\mathbb{Q},x)}.$
@@
[…]
@@
What we call the affine height density is simply called the natural density in the literature. The reason for our somewhat unusual terminology and notation will become clear in the next section when we have various subsets of projective spaces and their affine liftings. When $\mathcal{S}$ is a set of rational maps of degree $d$, it appears more natural to replace #$\#\mathbb{P}^{2d+1}(\mathbb{Q},x)$ by #$\#\mathcal{R}^{(d)}(x)$ in the definition of $\mathfrak{d}(S)$, $\underline{\mathfrak{d}}(S)$, and $\overline{\mathfrak{d}}(S)$. But this is equivalent to the given definition thanks to Proposition 2.6.
@@ 2.2. The box density and weak box densities
$\operatorname{Box}(\vec{r}):=\{(x_{1},\dots,x_{2d+2})\in\mathbb{Z}^{2d+2}:\ |x_{i}|\leq r_{i}\text{ for all }i\in[1,2d+2]\}.$
@@
$\mu(\mathcal{S}):=\lim_{r_{1},\dots,r_{2d+2}\rightarrow\infty}\frac{\#\mathcal{S}(\vec{r})}{\#\mathbb{P}^{2d+2}(\mathbb{Q},\vec{r})}$
[…]
@@
$\mu^{*}(S):=\lim_{r_{1},\dots,r_{2d+2}\rightarrow\infty}\frac{\#S(\vec{r})}{\#\operatorname{Box}(\vec{r})}$
[…]
@@
$\lim_{r_{1},\dots,\hat{r}_{n},\dots,r_{2d+2}\rightarrow\infty}:=\lim_{r_{1},\dots,r_{n-1},r_{n+1},\dots,r_{2d+2}\rightarrow\infty},$
@@
$\begin{split}&\underline{\mu_{n}}(\mathcal{S}):=\liminf_{r_{1},\dots,\hat{r}_{n},\dots,r_{2d+2}\rightarrow\infty}\left\{\liminf_{r_{n}\rightarrow\infty}\left(\frac{\#\mathcal{S}(\vec{r})}{\#\mathbb{P}^{2d+1}(\mathbb{Q},\vec{r})}\right)\right\}\ \text{and}\\ &\overline{\mu_{n}}(\mathcal{S}):=\limsup_{r_{1},\dots,\hat{r}_{n},\dots,r_{2d+2}\rightarrow\infty}\left\{\limsup_{r_{n}\rightarrow\infty}\left(\frac{\#\mathcal{S}(\vec{r})}{\#\mathbb{P}^{2d+1}(\mathbb{Q},\vec{r})}\right)\right\}.\end{split}$
@@
Let $S$ be a subset of $\mathbb{Z}^{2d+2}$, we define the lower affine $n$-th weak box density $\underline{\mu_{n}}^{*}(S)$, the upper affine $n$-th weak box density $\overline{\mu_{n}}^{*}(S)$, and the affine $n$-th weak box density $\mu_{n}^{*}(S)$ in a similar manner by replacing #$\#\mathbb{P}^{2d+1}(\mathbb{Q},\vec{r})$ by #$\#\operatorname{Box}(\vec{r})$.
We note that the definition of the affine densities in this subsection are taken from Poonen’s paper [Poo03]; however his terminologies do not include the word “affine” and his notations do not include the star superscript. Since $\mathcal{R}^{(d)}$ is naturally a subset of $\mathbb{P}^{2d+1}(\mathbb{Q})$, we reserve the simpler terminologies and notations for the projective setting. As in Remark 2.8, when $\mathcal{S}\subseteq\mathcal{R}^{(d)}$, it is natural to define the various densities of $\mathcal{S}$ using #$\#\mathcal{R}^{(d)}(\vec{r})$ instead of #$\#\mathbb{P}^{2d+1}(\mathbb{Q},\vec{r})$. This is equivalent to the given definition thanks to the box version of Proposition 2.4 and Proposition 2.6 established in the next section. For $\mathcal{S}\subseteq\mathbb{P}^{2d+1}(\mathbb{Q})$ (respectively $S\subseteq\mathbb{Z}^{2d+2}$), if the box density $\mu(\mathcal{S})$ (respectively $\mu^{*}(S)$) exists, then the height density $\mathfrak{d}(\mathcal{S})$ (respectively $\mathfrak{d}^{*}(S)$) and the weak box density $\mu_{n}(\mathcal{S})$ (respectively $\mu_{n}^{*}(S)$) exist and have the same value.

### v3

@@ 1.1. Background and motivation
For an elliptic curve defined over a number field, one can define the local minimal discriminant at each non-archimedean place and define the (global) minimal discriminant as the product of the local minimal discriminants [silvermanEC, Chapters VII–VIII]. Similarly, for a rational map defined over a number field, one can define the local minimal resultant at each non-archimedean place and define the (global) minimal resultant as their product, see [silvermantextbook, Chapter 4.11]. After homogenization, a rational map $\phi\in\mathbb{Q}(z)$ is given by a pair of homogeneous polynomials $F(X,Y),G(X,Y)\in\mathbb{Z}[X,Y]$ without common factors in $\mathbb{Z}[X,Y]$. We say that $\phi$ is globally minimal or has globally minimal resultant if the resultant of $F$ and $G$ is equal to the minimal resultant of $\phi$. Over a complete algebraically closed non-archimedean valued field, the notion of (local) minimal resultants give rise to the so called minimal resultant locus which has interesting analytic and moduli-theoretic characterizations [Rumely2015, Rumely2017]. Over $\mathbb{Q}$, globally minimal maps play an important role in a conjecture of Silverman [silvermantextbook, Conjecture 3.47] on the uniformity of number of integral points in orbits. This is, in turn, analogous to a conjecture of Lang [Lang] for integral points on elliptic curves after work of Dem’janenko.
In recent years, there have been various results concerning statistical aspects of the arithmetic dynamics on $\mathbb{P}^{1}$, for example [HH2019, H2021a, H2021b, DHHJ2024]. However, the nature of these results involves a stochastic process in which one successively composes a random rational map from a finite set endowed with a probability measure. On the other hand, our work continues the classical analogy program between arithmetic dynamics on $\mathbb{P}^{1}$ and elliptic curves. In arithmetic statistics of elliptic curves, considerable interest lies in questions regarding the distribution of sets $\mathcal{S}$ with certain properties. For instance, the rank distribution conjecture of Katz and Sarnak predicts that exactly $1/2$ (resp. $1/2$) of the elliptic curves over $\mathbb{Q}$ have rank 0 (resp. $1$). Cremona and Sadek [cremonasadek] consider the density of global Weierstrass models of elliptic curves over $\mathbb{Q}$ with prescribed reduction types at the primes. We prove analogues of results of Cremona and Sadek for models for dynamical systems, arising from rational maps of prescribed degree $d\geq 2$ in $\mathbb{P}^{1}$ with coefficients in $\mathbb{Q}$.
@@ 1.2. Main results
$P(\phi)\mathrel{\mathop{\ordinarycolon}}=[A_{0}\mathrel{\mathop{\ordinarycolon}}\cdots\mathrel{\mathop{\ordinarycolon}}A_{d}\mathrel{\mathop{\ordinarycolon}}B_{0}\mathrel{\mathop{\ordinarycolon}}\cdots\mathrel{\mathop{\ordinarycolon}}B_{d}]\in\mathbb{P}^{2d+1}(\mathbb{Q}).$
The height of $\phi$, denoted $H(\phi)$, is defined to be the absolute multiplicative Weil height of the point $P(\phi)$ [BomGub, Chapter 1]. For any $x>0$, let $\mathcal{R}^{(d)}(x)=\{\phi\in\mathcal{R}^{(d)}\mathrel{\mathop{\ordinarycolon}}\ H(\phi)\leq x\}$. It is seen that
$\mathcal{R}^{(d)}(x)=\frac{(2x)^{2d+2}}{2\zeta(2d+2)}+O\left(x^{2d+1}\log x\right),$
@@
$\liminf_{x\rightarrow\infty}\left(\frac{\#\{\phi\in\mathcal{R}^{(d)}(x)\mid\phi\in\mathcal{S}\}}{\#\mathcal{R}^{(d)}(x)}\right)>\frac{c}{100}.$
Inspired by Poonen’s work [Poonensquarefree], there is a more general density to consider. First, it is possible to regard $\mathcal{R}^{(d)}\subset\mathbb{P}^{1}(\mathbb{Q})$ as subsets of $\mathbb{Z}^{2d+2}$, see Notation 2.2. Then one may consider the box density in [Poonensquarefree, p. 355]. There are weaker versions of this called the weak box densities, cf. Theorem 3.2 of loc. cit. More specifically, for $\mathcal{S}\subset\mathcal{R}^{(d)}$ and for each natural number $n\in[1,2d+2]$, we have the notion of the weak box density $\mu_{n}(\mathcal{S})$, see Definition 2.11.
@@
The following analogue of the result of Cremona-Sadek [cremonasadek] arises: with respect to the box density, does a positive proportion of rational maps have globally minimal resultant? This question, along with related topics, will be explored in a future work.
@@ 1.3. Methodology
We briefly discuss the techniques used in the proof of Theorems A, B and C. A classical result of Schanuel [schanuel] gives the asymptotic count for the number of rational points in projective space with height $\leq x$. For two generic homogeneous polynomials $F(X,Y)$ and $G(X,Y)$ of degree $d$, their resultant is a homogeneous polynomial in the coefficients of $F$ and $G$. Hence $F$ and $G$ have a common factor if and only if the resultant polynomial vanishes. A result by Schwartz and Zippel (cf. Lemma 2.5) provides an upper bound on the number of zeros of this polynomial with height $\leq x$. Combining this result with Schanuel’s result implies that
$\mathcal{R}^{(d)}(x)\sim\frac{(2x)^{2d+2}}{2\zeta(2d+2)},$
as detailed in Proposition 2.6. Following Poonen’s work [Poonensquarefree], we can also study the “box version” of the above results.
In Section 3, we study properties of subsets $\mathcal{S}\subseteq\mathcal{R}^{(d)}$ defined by the so called local conditions. We define local densities at each prime $p$. Proposition 3.4 employs a sieve-theoretic argument to show that, if a certain admissibility condition is met, the density of $\mathcal{S}$ is the product of local densities at all primes $p$. These results are applied in Section 4 to prove Theorems A and B. Although these local conditions are insufficient to describe the entire collection of $\phi$ with minimal resultant, they provide positive lower bounds on the density. This relies on explicit local density calculations, with a main innovation involving counting the number of $\mathbb{Z}/p^{2}\mathbb{Z}$-points on the integral scheme defined by the resultant polynomial and its partial derivatives. Some similar arguments have been employed in [Poonensquarefree]. In Section 5, an explicit primary decomposition for this integral scheme is used to obtain an explicit lower bound on the weak box densities when $d=2$. Our computations are aided by Macaulay2.
@@ 1.4. Acknowledgements
Data availability statement
All relevant data has been included in the paper.
@@ 2. The height density, box density, and weak box densities
We define the height $H(\phi)$ of $\phi\in\mathcal{R}^{(d)}$ as the absolute multiplicative Weil height of the corresponding point in $\mathbb{P}^{2d+1}(\mathbb{Q})$. Then we define the height density of a subset $\mathcal{S}$ of $\mathcal{R}^{(d)}$. Using classical results of Schanuel [schanuel] for the distribution of rational points on projective spaces and the Schwartz-Zippel lemma [Schwartz, Zippel], we derive an asymptotic formula for
$\mathcal{R}^{(d)}(x)\mathrel{\mathop{\ordinarycolon}}=\{\phi\in\mathcal{R}^{(d)}\mathrel{\mathop{\ordinarycolon}}\ H(\phi)\leq x\}.$
We also recall the notion of box density and weak box densities for subsets of $\mathbb{Z}^{2d+2}$ from Poonen’s work [Poonensquarefree]. We have the analogues of Schanuel’s result and an asymptotic formula for counting $\mathcal{R}^{(d)}$ in a box; however the statement of these analogues is delayed until the next section since the proof relies on certain sieving results. Finally we define the box density and weak box densities for a subset $\mathcal{S}$ of $\mathcal{R}^{(d)}$.
@@ 2.1. The height density
$P=P(\phi)\mathrel{\mathop{\ordinarycolon}}=[A_{0}\mathrel{\mathop{\ordinarycolon}}\cdots\mathrel{\mathop{\ordinarycolon}}A_{d}\mathrel{\mathop{\ordinarycolon}}B_{0}\mathrel{\mathop{\ordinarycolon}}\cdots\mathrel{\mathop{\ordinarycolon}}B_{d}]\in\mathbb{P}^{2d+1}(\mathbb{Q}).$
@@
(2.1) $H(\phi)=H(P)\mathrel{\mathop{\ordinarycolon}}=\prod_{v}\max\{|A_{i}|_{v},|B_{j}|_{v}\mathrel{\mathop{\ordinarycolon}}\ 0\leq i,j\leq d\},$
@@
Thus we count models for rational maps, and not isomorphism classes. Nonetheless, this seems to be natural way to order rational maps, the definition is analogous to counting Weierstrass equations for elliptic curves according to their height (cf. [cremonasadek]). For a detailed treatment of heights, we refer to [BomGub].
@@
We shall also denote the coordinates of $Q$ by $(A_{0},\dots,A_{d},B_{0},\dots,B_{d})$, i.e., $A_{i}\mathrel{\mathop{\ordinarycolon}}=C_{i+1}$ and $B_{j}\mathrel{\mathop{\ordinarycolon}}=C_{d+j+2}$. Note that if $Q$ satisfies (2), then exactly one of the points $Q$ or Q-Q satisfies (1).
@@
$H(\phi)=\max\{|A_{i}|,|B_{j}|\mathrel{\mathop{\ordinarycolon}}0\leq i,j\leq d\},$
see [BomGub, Chapter 1].
@@
$\mathcal{S}(x)=\{P\in\mathcal{S}\mathrel{\mathop{\ordinarycolon}}\ H(P)\leq x\}.$
@@
(2.2) $\mathbb{P}^{2d+1}(\mathbb{Q},x)=\frac{1}{2\zeta(2d+2)}(2x)^{2d+2}+O\left(x^{2d+1}\log x\right).$
@@
This is from [schanuel, Theorem 1]. ∎
@@
$\{(x_{1},\ldots,x_{n})\in I_{1}\times\cdots\times I_{n}\mathrel{\mathop{\ordinarycolon}}\ f(x_{1},\ldots,x_{n})=0\}$
@@
This was proven independently by Schwartz [Schwartz] and Zippel [Zippel]. ∎
@@
$\mathcal{R}^{(d)}(x)=\frac{1}{2\zeta(2d+2)}(2x)^{2d+2}+O\left(x^{2d+1}\log x\right),$
@@
$(\mathbb{P}^{2d+1}(\mathbb{Q},x)\setminus\mathcal{R}^{(d)}(x))=O(x^{2d+1})$
@@
$\mathfrak{d}(\mathcal{S})\mathrel{\mathop{\ordinarycolon}}=\lim_{x\rightarrow\infty}\frac{\#\mathcal{S}(x)}{\#\mathbb{P}^{2d+1}(\mathbb{Q},x)},$
@@
$\underline{\mathfrak{d}}(\mathcal{S})\mathrel{\mathop{\ordinarycolon}}=\liminf_{x\rightarrow\infty}\frac{\#\mathcal{S}(x)}{\#\mathbb{P}^{2d+1}(\mathbb{Q},x)}\text{ and }\overline{\mathfrak{d}}(\mathcal{S})\mathrel{\mathop{\ordinarycolon}}=\limsup_{x\rightarrow\infty}\frac{\#\mathcal{S}(x)}{\#\mathbb{P}^{2d+1}(\mathbb{Q},x)}.$
@@
provided the limit exists. The lower and upper affine height densities $\underline{\mathfrak{d}}^{*}$ and $\overline{\mathfrak{d}}^{*}$ are defined using $\liminf$ and $\limsup$ respectively.
@@
What we call the affine height density is simply called the natural density in the literature. The reason for our somewhat unusual terminology and notation will become clear in the next section when we have various subsets of projective spaces and their affine liftings. When $\mathcal{S}$ is a set of rational maps of degree $d$, it appears more natural to replace $\mathbb{P}^{2d+1}(\mathbb{Q},x)$ by $\mathcal{R}^{(d)}(x)$ in the definition of $\mathfrak{d}(S)$, $\underline{\mathfrak{d}}(S)$, and $\overline{\mathfrak{d}}(S)$. But this is equivalent to the given definition thanks to Proposition 2.6.
@@ 2.2. The box density and weak box densities
$\operatorname{Box}(\vec{r})\mathrel{\mathop{\ordinarycolon}}=\{(x_{1},\dots,x_{2d+2})\in\mathbb{Z}^{2d+2}\mathrel{\mathop{\ordinarycolon}}\ |x_{i}|\leq r_{i}\text{ for all }i\in[1,2d+2]\}.$
@@
$\mu(\mathcal{S})\mathrel{\mathop{\ordinarycolon}}=\lim_{r_{1},\dots,r_{2d+2}\rightarrow\infty}\frac{\#\mathcal{S}(\vec{r})}{\#\mathbb{P}^{2d+2}(\mathbb{Q},\vec{r})}$
provided the limit exists. The lower and upper box densities $\underline{\mu}(\mathcal{S})$ and $\overline{\mu}(\mathcal{S})$ are defined using $\liminf$ and $\limsup$ respectively.
@@
$\mu^{*}(S)\mathrel{\mathop{\ordinarycolon}}=\lim_{r_{1},\dots,r_{2d+2}\rightarrow\infty}\frac{\#S(\vec{r})}{\#\operatorname{Box}(\vec{r})}$
provided the limit exists. The lower and upper affine box densities $\underline{\mu}^{*}(S)$ and $\overline{\mu}^{*}(S)$ are defined using $\liminf$ and $\limsup$ respectively.
@@
$\lim_{r_{1},\dots,\hat{r}_{n},\dots,r_{2d+2}\rightarrow\infty}\mathrel{\mathop{\ordinarycolon}}=\lim_{r_{1},\dots,r_{n-1},r_{n+1},\dots,r_{2d+2}\rightarrow\infty},$
@@
$\begin{split}&\underline{\mu_{n}}(\mathcal{S})\mathrel{\mathop{\ordinarycolon}}=\liminf_{r_{1},\dots,\hat{r}_{n},\dots,r_{2d+2}\rightarrow\infty}\left\{\liminf_{r_{n}\rightarrow\infty}\left(\frac{\#\mathcal{S}(\vec{r})}{\#\mathbb{P}^{2d+1}(\mathbb{Q},\vec{r})}\right)\right\}\ \text{and}\\ &\overline{\mu_{n}}(\mathcal{S})\mathrel{\mathop{\ordinarycolon}}=\limsup_{r_{1},\dots,\hat{r}_{n},\dots,r_{2d+2}\rightarrow\infty}\left\{\limsup_{r_{n}\rightarrow\infty}\left(\frac{\#\mathcal{S}(\vec{r})}{\#\mathbb{P}^{2d+1}(\mathbb{Q},\vec{r})}\right)\right\}.\end{split}$
@@
Let $S$ be a subset of $\mathbb{Z}^{2d+2}$, we define the lower affine $n$-th weak box density $\underline{\mu_{n}}^{*}(S)$, the upper affine $n$-th weak box density $\overline{\mu_{n}}^{*}(S)$, and the affine $n$-th weak box density $\mu_{n}^{*}(S)$ in a similar manner by replacing $\mathbb{P}^{2d+1}(\mathbb{Q},\vec{r})$ by $\operatorname{Box}(\vec{r})$.
We note that the definition of the affine densities in this subsection are taken from Poonen’s paper [Poonensquarefree]; however his terminologies do not include the word “affine” and his notations do not include the star superscript. Since $\mathcal{R}^{(d)}$ is naturally a subset of $\mathbb{P}^{2d+1}(\mathbb{Q})$, we reserve the simpler terminologies and notations for the projective setting. As in Remark 2.8, when $\mathcal{S}\subseteq\mathcal{R}^{(d)}$, it is natural to define the various densities of $\mathcal{S}$ using $\mathcal{R}^{(d)}(\vec{r})$ instead of $\mathbb{P}^{2d+1}(\mathbb{Q},\vec{r})$. This is equivalent to the given definition thanks to the box version of Proposition 2.4 and Proposition 2.6 established in the next section. For $\mathcal{S}\subseteq\mathbb{P}^{2d+1}(\mathbb{Q})$ (respectively $S\subseteq\mathbb{Z}^{2d+2}$), if the box density $\mu(\mathcal{S})$ (respectively $\mu^{*}(S)$) exists, then the height density $\mathfrak{d}(\mathcal{S})$ (respectively $\mathfrak{d}^{*}(S)$) and the weak box density $\mu_{n}(\mathcal{S})$ (respectively $\mu_{n}^{*}(S)$) exist and have the same value.